# Mixture of Experts-augmented Deep Unfolding for Activity Detection in IRS-aided Systems

### v1

@@ V Conclusions
In this paper, we proposed a model-driven deep learning approach for activity detection in IRS-aided systems with mixed types of channel fading. Specifically, we derived an approximated covariance-based formulation and introduced a deep unfolding network based on the projected gradient method. By further incorporating a MoE network, the proposed method does not require prior knowledge of which device belongs to which channel type. Simulation results showed that the MoE-augmented model-driven deep unfolding method achieves better detection performance than traditional covariance-based methods and black-box neural network designs. Furthermore, the performance loss from not knowing the fading channel type of each device is minimal compared to the unfolded network with perfect information.
Appendix A Proof of Proposition 1
Let $\mathbf{h}_{k}^{\text{NLoS}}\sim\mathcal{CN}\left(\mathbf{0},\mathbf{I}_{N}\right)$ and $\mathbf{g}_{m}^{\text{NLoS}}\sim\mathcal{CN}\left(\mathbf{0},\mathbf{I}_{N}\right)$ be the non-line-of-site component of $\mathbf{h}_{k}$ and $\mathbf{g}_{m}$, respectively. Based on [16, Proposition 1], we have $(\mathbf{h}_{k}^{\text{NLoS}})^{H}\mathbf{\Theta}\mathbf{g}_{m}^{\text{NLoS}}\sim\mathcal{CN}\left(\mathbf{0},N\mathbf{I}_{M}\right)$ when $N$ is sufficiently large. Therefore, $\mathbf{Y}$ is a complex Gaussian distributed matrix. The mean of $\mathbf{y}_{m}$ is given by
$\mathbb{E}\left[\mathbf{y}_{m}\right]&=\sum_{k\in\mathcal{K}_{1}}{a_{k}}\mathbf{s}_{k}(\sqrt{\frac{\kappa_{\text{U}}}{1+\kappa_{\text{U}}}}\mathbf{h}_{k}^{\text{LoS}})^{H}\mathbf{\Theta}\sqrt{\frac{\kappa_{\text{B}}}{1+\kappa_{\text{B}}}}\mathbf{g}_{m}^{\text{LoS}}\\ &+\sum_{k\in\mathcal{K}_{2}}{a_{k}}\mathbf{s}_{k}\sqrt{\frac{\kappa_{\text{U}}}{1+\kappa_{\text{U}}}}\mathbf{f}_{k}^{\text{LoS}}(m)^{*}\end{aligned}$
and
$\mathbb{E}\left[(\mathbf{y}_{m}-\mathbb{E}\left[\mathbf{y}_{m}\right])(\mathbf{y}_{m}-\mathbb{E}\left[\mathbf{y}_{m}\right])^{H}\right]$
$=$ $\!\sum_{k\in\mathcal{K}_{1}}{a_{k}}\Xi_{m,k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}\!+\!\sum_{k\in\mathcal{K}_{2}}\frac{a_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}}{1+\kappa_{\text{U}}}+\!\sum_{k\in\mathcal{K}_{3}}{a_{k}}\mathbf{s}_{k}\mathbf{s}_{k}^{H}$
$+$ $\sigma_{w}^{2}\mathbf{I}_{M},$
where $\Xi_{m,k}=\frac{N+\kappa_{\text{B}}\|{\mathbf{g}}_{m}^{\text{LoS}}\|^{2}+\kappa_{\text{U}}\|{\mathbf{h}}_{k}^{\text{LoS}}\|^{2}}{(1+\kappa_{\text{B}})(1+\kappa_{\text{U}})}$. Once obtaining $\bar{\mathbf{y}}_{m}\triangleq\mathbb{E}\left[\mathbf{y}_{m}\right]$ and $\mathbf{\Sigma}_{m}\triangleq\mathbb{E}\left[(\mathbf{y}_{m}-\mathbb{E}\left[\mathbf{y}_{m}\right])\right.\left.(\mathbf{y}_{m}-\mathbb{E}\left[\mathbf{y}_{m}\right])^{H}\right]$, we have (1).

### v2

@@ V Conclusions
In this paper, we proposed a model-driven deep learning approach for activity detection in IRS-aided systems with mixed types of channel fading. Specifically, we derived an approximated covariance-based formulation and introduced a deep unfolding network based on the projected gradient method. By further incorporating a MoE network, the proposed method does not require prior knowledge of which device belongs to which channel type. Simulation results showed that the MoE-augmented model-driven deep unfolding method achieves better detection performance than traditional covariance-based methods and black-box neural network designs. Furthermore, the performance loss from not knowing the fading channel type of each device is minimal compared to the unfolded network with perfect channel type information.